OpenStudy (anonymous):

find three solutions of the equation, y = 3x + 4. a. (-2 , -2) , (1 , 7) b. (0 , 4) , ( 3, 11) , (-2 , -2) c. (-2 , -2) , ( 1 , 7) , ( 0 , 4) d. ( 1, 7) , ( 0, 4 ) , ( -1 , 0)

OpenStudy (jack1):

....? A has only 2 solutions...?

OpenStudy (anonymous):

so its a ?

OpenStudy (jack1):

im saying it cant be a or a is missing something...?

OpenStudy (anonymous):

help i have 2 mins left just give me any answer please!

OpenStudy (jack1):

try each y = 3x + 4, so if x = 0 and y = 4 4 = 3(0) + 4 4 = 0 +4 4 = 4 LHS = RHS, so could be B, test the other points to confirm

OpenStudy (jack1):

your turn: y = 3x + 4, so if x = 3 and y = 11 solve and see if Left Hand Side = Right Hand Side

OpenStudy (anonymous):

is it b ? c ? a ?

OpenStudy (anonymous):

30 sec left

OpenStudy (jack1):

im not doing your test for u... u solve

OpenStudy (mokeira):

(a) and (c) just substitute the values